# Tapor: 3D Hand Pose Reconstruction with Fully Passive Thermal Sensing for Around-device Interactions

### v1

@@ 7.2. Thermal Array-based Sensing
Thermal array sensors are gaining prominence due to their versatility in a variety of sensing applications. These applications include human detection (Cerutti et al., 2019; Tanaka et al., 2014), occupancy estimation (Sahoo and Lone, 2023; Chidurala and Li, 2021; Tyndall et al., 2016), daily activity recognition (Muthukumar et al., 2022; Rezaei et al., 2022; Polla et al., 2019), fall detection (He et al., 2023; Naser et al., 2022; Zhong et al., 2021; Rezaei et al., 2021), and indoor localization (Bouazizi et al., 2022; Faulkner et al., 2021; Hevesi et al., 2014). For gesture recognition, Tateno et al.(Tateno et al., 2019) developed an in-car interface system using a thermal array sensor and a CNN model. (Vandersteegen et al., 2020) proposed a low-latency system using a temporal convolution network for thermal array map sequences. However, these systems are limited to recognizing predefined gestures and require retraining to support new gestures. In contrast, Tapor enables 3D hand pose reconstruction, supporting more flexible and scalable gesture recognition. A relevant study, FingerTrak (Hu et al., 2020), uses multiple thermal array sensors in a wristband for hand pose tracking, leading to a costly intrusive design. Conversely, Tapor leverages a single thermal array sensor for non-intrusive 3D hand pose reconstruction, streamlining hardware and lowering costs for wider applications.

### v2

@@ 7.2. Thermal Array-based Sensing
Thermal array sensors are gaining prominence due to their versatility in a variety of sensing applications. These applications include human detection (Cerutti et al., 2019; Tanaka et al., 2014), occupancy estimation (Sahoo and Lone, 2023; Chidurala and Li, 2021; Tyndall et al., 2016), daily activity recognition (Muthukumar et al., 2022; Rezaei et al., 2022; Polla et al., 2019), fall detection (He et al., 2023; Naser et al., 2022; Zhong et al., 2021; Rezaei et al., 2021), and indoor localization (Bouazizi et al., 2022; Faulkner et al., 2021; Hevesi et al., 2014). For gesture recognition, although there has been extensive work using thermal cameras, such as Birkeland et al.(Birkeland et al., 2024), who provided a large thermal hand gesture dataset captured with the FLIR Lepton thermal camera, and ThermoHands (Ding et al., 2024), which introduced an egocentric benchmark dataset and a transformer-based baseline for robust 3D hand pose estimation under diverse real-world conditions, research on thermal array sensors remains limited.
As a new and emerging sensing modality, thermal arrays offer unique advantages such as extremely low resolution, low signal-to-noise ratio, low power consumption, and significantly lower cost compared to thermal cameras. Despite these promising features, only a few studies have explored their potential for gesture recognition or hand pose estimation. Notable efforts in this direction include Tateno et al.(Tateno et al., 2019), who developed an in-car interface system using a thermal array sensor and a CNN model, and Vandersteegen et al.(Vandersteegen et al., 2020), who proposed a low-latency gesture recognition system based on temporal convolutional networks applied to thermal array map sequences. A particularly relevant study, FingerTrak (Hu et al., 2020), employs multiple thermal array sensors embedded in a wristband to continuously track hand poses. While this design effectively addresses occlusion issues, it introduces significant hardware complexity and imposes wearability constraints. In contrast, TAPOR is tailored for ambient, non-wearable scenarios. It leverages a single low-cost thermal array sensor to enable seamless around-device interaction, significantly simplifying deployment and enhancing usability.
TAPOR distinguishes itself by enabling 3D hand pose reconstruction solely from a thermal array sensor, offering a scalable, privacy-preserving, and cost-effective solution well-suited for widespread IoT applications.
@@ 8. Conclusions
Acknowledgements.
This work is supported by the NSFC under grant No. 62222216 and Hong Kong RGC ECS under grant No. 27204522.